# Winning Criteria for Open Games: A Game-Theoretic Approach to Prefix Codes

### v1

@@ Abstract
We study two-player games with alternating moves played on infinite trees. Our main focus is on the case where the trees are full (regular) and the winning set is open (with respect to the product topology on the tree). Gale and Stewart showed that in this setting one of the players always has a winning strategy, though it is not known in advance which player. We present simple necessary conditions for the first player to have a winning strategy, and establish an equivalence between winning sets that guarantee a win for the first player and maximal prefix codes. Using this equivalence, we derive a necessary algebraic condition for winning, and exhibit a family of games for which this algebraic condition is in fact equivalent to winning. We introduce the concept of coverings, and show that by covering the graph with an infinite labeled tree corresponding to the free group, we can derive a simple trait of maximal prefix codes.
@@ 1 Introduction
Section 4 introduces Nielsen–Schreier theory, while the resulting algebraic conditions are established in Section 5. Finally, the notion of covering is introduced in Section 6.
@@ 2.3 Redefining Open Winning Set: Hausdorff Dimension Reduction
Let $T=\mathcal{A}^{\leq\mathbb{N}}$ and let $W=\bigcup_{p\in Z}[T_{p}]$ be an open set. If $\dim_{H}(Z_{\mathrm{concat}})<\tfrac{1}{2}$, then
$\sum_{p\in Z}\frac{1}{|\mathcal{A}|^{\lfloor\mathrm{len}(p)/2\rfloor}}<1.$
Proof. Since $\dim_{H}(Z_{\mathrm{concat}})<\tfrac{1}{2}$, the definition of Hausdorff dimension implies that for any cover $\{U_{i}\}_{i\in I}$ of $W_{\mathrm{concat}}$ by basic open sets of diameter $|\mathcal{A}|^{-2n}$, we have
$\sum_{i\in I}\left(|\mathcal{A}|^{-2n}\right)^{1/2}\;\longrightarrow\;0\qquad\text{as }n\to\infty.$
Equivalently, letting
$\mathrm{Cov}(W,2n):=\{\,p\in T:\mathrm{len}(p)=2n\text{ and }p\in W_{\mathrm{concat}}\,\},$
we obtain
$|\mathrm{Cov}(W,2n)|\cdot|\mathcal{A}|^{-n}\;\xrightarrow[n\to\infty]{}\;0.$
@@
Now fix $p\in Z$. We replace $p$ by a uniformly expanded family of extensions of length $2n$, constructed by inserting a fixed symbol $a_{0}\in\mathcal{A}$ in every even position and allowing all symbols of $\mathcal{A}$ to appear in the intervening odd positions.
If $p$ has even length, define
@@
$\mathrm{Extend}_{n}(p)=\Bigl\{p\circ\langle a_{0},a_{1}\rangle\circ\langle a_{0},a_{2}\rangle\circ\cdots\circ\langle a_{0},a_{n-\lfloor\mathrm{len}(p)/2\rfloor}\rangle:\;a_{1},\ldots,a_{n-\lfloor\mathrm{len}(p)/2\rfloor}\in\mathcal{A}\Bigr\}.$
If $p$ has odd length, define
$\mathrm{Extend}_{n}(p)=\Bigl\{p\circ\langle a_{1}\rangle\circ\langle a_{0},a_{2}\rangle\circ\cdots\circ\langle a_{0},a_{n-\lfloor\mathrm{len}(p)/2\rfloor}\rangle:\;a_{1},\ldots,a_{n-\lfloor\mathrm{len}(p)/2\rfloor}\in\mathcal{A}\Bigr\}.$
Every $q\in\mathrm{Extend}_{n}(p)$ satisfies $\lfloor\mathrm{len}(q)/2\rfloor=n$, and
$|\mathrm{Extend}_{n}(p)|=|\mathcal{A}|^{\,n-\lfloor\mathrm{len}(p)/2\rfloor}.$
Thus the contribution of $p$ to the sum
$\sum_{q\in\mathrm{Extend}_{n}(p)}\frac{1}{|\mathcal{A}|^{\lfloor\mathrm{len}(q)/2\rfloor}}$
is
$\sum_{q\in\mathrm{Extend}_{n}(p)}\frac{1}{|\mathcal{A}|^{n}}=|\mathcal{A}|^{\,n-\lfloor\mathrm{len}(p)/2\rfloor}\cdot\frac{1}{|\mathcal{A}|^{n}}=\frac{1}{|\mathcal{A}|^{\lfloor\mathrm{len}(p)/2\rfloor}}.$
Therefore, the original sum can be rewritten as
$\sum_{p\in Z}\frac{1}{|\mathcal{A}|^{\lfloor\mathrm{len}(p)/2\rfloor}}=\sum_{p\in Z}\sum_{q\in\mathrm{Extend}_{n}(p)}\frac{1}{|\mathcal{A}|^{n}}=\sum_{q\in\mathrm{Extend}_{n}(Z)}\frac{1}{|\mathcal{A}|^{n}}.$
But the set $\mathrm{Extend}_{n}(Z)$ consists exactly of the prefixes in $\mathrm{Cov}(W,2n)$, and hence
$\sum_{p\in Z}\frac{1}{|\mathcal{A}|^{\lfloor\mathrm{len}(p)/2\rfloor}}=|\mathrm{Cov}(W,2n)|\cdot|\mathcal{A}|^{-n}.$
For sufficiently large $n$, the right-hand side is strictly less than $1$, completing the proof.
@@ 4 Introduction to Nielsen-Schreier Theory
In Figure 1, we illustrate the Schreier graphs of two free subgroups of $F(\mathcal{A})$ with $\mathcal{A}=\{a,b\}$, one of which being the trivial subgroup, denoted $\langle\rangle$, containing only the empty word [^6] The Schreier graph of the trivial subgroup is better knows as the Cayley graph of the free group..
@@ 5 Algebraic Criteria Using Maximal Prefix Codes
Notice that the proof of the previous Theorem, showed that in fact, we have at least $|\{p\in Z:\mathrm{len}(p)\text{ is maximal}\}|/|\mathcal{A}|\geq 1$ different ways to choose $Z^{\prime}$, as the initial choice of the maximal length word was arbitrary.

### v2

@@ Abstract
We study two-player games with alternating moves played on infinite trees. Our main focus is on the case where the trees are full (regular) and the winning set is open (with respect to the product topology on the tree). Gale and Stewart showed that in this setting one of the players always has a winning strategy, though it is not known in advance which player. We present simple necessary conditions for the first player to have a winning strategy, and establish an equivalence between winning sets that guarantee a win for the first player and maximal prefix codes. Using this equivalence, we derive a necessary algebraic condition for winning, and exhibit a family of games for which this algebraic condition is in fact equivalent to winning. We introduce the concept of coverings, and show that by covering the tree of the game with an infinite labeled tree corresponding to the free group, we can use “game-theoretic tools” to derive a simple trait of maximal prefix codes.
@@ 1 Introduction
At Section 7 we discuss some extension of the results to games in which the action set of each player is countable.
@@
Section 4 introduces Nielsen–Schreier theory, while the resulting algebraic conditions are established in Section 5. The notion of covering is introduced in Section 6, and finally, further discussion about games played with a countable alphabet are presented in Section 7.
@@ 2.3 Redefining Open Winning Set: Hausdorff Dimension Reduction
Let $T=\mathcal{A}^{\leq\mathbb{N}}$ and let $W=\bigcup_{p\in Z}[T_{p}]$ be an open set, with $Z\subseteq T$ being a finite collection of even positions. Then $\dim_{H}(Z_{\mathrm{concat}})<\tfrac{1}{2}$, if and only if
$\sum_{p\in Z}\frac{1}{|\mathcal{A}|^{\mathrm{len}(p)/2}}<1.$
Before giving the proof, we state a result that will be used below. This statement is a corollary of the Hausdorff dimension game introduced by Das, Fishman, Simmons, and Urbański [11]; see also [5] for an exposition in the dyadic setting.
Theorem 2.19
Let $T=\mathcal{A}^{\leq\mathbb{N}}$ and let $S\subseteq[T]$ be a Borel set. If there exist integers $k\geq 1$ and a set $A\subseteq\mathcal{A}^{k}$ with $|A|\geq|\mathcal{A}|^{kd}$ and $A_{\mathrm{concat}}\subseteq S$, then $\dim_{H}(S)\geq d$.
Proof of Lemma 2.18. We recall the definition of Hausdorff dimension.
Let $\delta>0$. A $\delta$-cover of $Z_{\mathrm{concat}}$ is a countable collection $\{U_{i}\}_{i\in I}$ such that $Z_{\mathrm{concat}}\subseteq\bigcup_{i\in I}U_{i}$ and $\operatorname{diam}(U_{i})<\delta$ for every $i\in I$. The $d$-dimensional Hausdorff measure of $Z_{\mathrm{concat}}$ is
$H^{d}(Z_{\mathrm{concat}})=\liminf_{\delta\to 0}\left\{\sum_{i\in I}(\operatorname{diam}U_{i})^{d}\;\middle|\;\{U_{i}\}_{i\in I}\text{ is a }\delta\text{-cover of }Z_{\mathrm{concat}}\right\},$
and the Hausdorff dimension is
$\dim_{H}(Z_{\mathrm{concat}})=\inf\{d\geq 0:\;H^{d}(Z_{\mathrm{concat}})=0\}.$
@@
Let $2n=\max_{p\in Z}\operatorname{len}(p)$, and fix $p\in Z$. We replace $p$ by a uniformly expanded family of extensions of length $2n$, constructed by inserting a fixed symbol $a_{0}\in\mathcal{A}$ in every even position and allowing every symbol of $\mathcal{A}$ to appear in the intervening odd positions.
Recall that we have assumed every $p\in Z$ has even length. Define
@@
$\operatorname{Extend}_{n}(p):=\Bigl\{p\circ\langle a_{0},a_{1}\rangle\circ\langle a_{0},a_{2}\rangle\circ\cdots\circ\langle a_{0},a_{\,n-(\operatorname{len}(p)/2)}\rangle:a_{1},\dots,a_{\,n-(\operatorname{len}(p)/2)}\in\mathcal{A}\Bigr\}.$
Every $q\in\operatorname{Extend}_{n}(p)$ satisfies $\operatorname{len}(q)=2n$, and
$\bigl|\operatorname{Extend}_{n}(p)\bigr|=|\mathcal{A}|^{\,n-(\operatorname{len}(p)/2)}.$
Set
$W_{n}:=\left(\bigcup_{p\in Z}\operatorname{Extend}_{n}(p)\right)_{\mathrm{concat}}.$
We claim that if $\dim_{H}(Z_{\mathrm{concat}})<1/2$ then $\dim_{H}(W_{n})<1/2$. To see this, let $\{U_{i}\}_{i\in I}$ be a $\delta$-cover of $Z_{\mathrm{concat}}$. We may assume each $U_{i}$ is a basic open set: if $\operatorname{diam}(U_{i})=0$ then $U_{i}$ contains a single point and can be replaced by a basic open set of arbitrarily small diameter; moreover,
$\operatorname{diam}(U_{i})=|\mathcal{A}|^{-\min\{k:\exists x,y\in U_{i},\;x_{k}\neq y_{k}\}},$
so replacing $U_{i}$ by a containing basic open set does not increase its diameter.
Since the infimum defining $H^{d}$ may be taken over basic open covers, we may further assume each $U_{i}$ has the form $T_{p}]$, where $p$ is a prefix of a concatenation $p_{1}\circ\cdots\circ p_{k}$ with $p_{1},\dots,p_{k}\in Z$. For such a basic open set $U_{i}=[T_{p}]$ with $p=p_{1}\circ\cdots\circ p_{k-1}\circ p^{\prime}$ and $p^{\prime}\preceq p_{k}\in Z$, define
$\operatorname{Extend}_{n}(U_{i}):=\bigl\{[T_{q}]:q=q_{1}\circ\cdots\circ q_{k-1}\circ p^{\prime},\;q_{j}\in\operatorname{Extend}_{n}(p_{j})\ \text{for }1\leq j\leq k-1\bigr\}.$
Then $\bigcup_{i\in I}\operatorname{Extend}_{n}(U_{i})$ is an open cover of $W_{n}$.
Because $\dim_{H}(Z_{\mathrm{concat}})<1/2$, there exists $\epsilon_{0}>0$ such that $H^{1/2-\epsilon_{0}}(Z_{\mathrm{concat}})=0$. Put $\epsilon=\epsilon_{0}/(2n)$. We claim that
$\sum_{i\in I}\sum_{U^{\prime}\in\operatorname{Extend}_{n}(U_{i})}(\operatorname{diam}U^{\prime})^{\,1/2-\epsilon}\leq\sum_{i\in I}(\operatorname{diam}U_{i})^{\,1/2-\epsilon_{0}}.$
To verify this inequality, fix $U_{i}=[T_{p}]$ with $p=p_{1}\circ\cdots\circ p_{k-1}\circ p^{\prime}$ as above. Then
$\sum_{U^{\prime}\in\operatorname{Extend}_{n}(U_{i})}(\operatorname{diam}U^{\prime})^{\,1/2-\epsilon}=\sum_{\begin{subarray}{c}q_{j}\in\operatorname{Extend}_{n}(p_{j})\\ j=1,\dots,k-1\end{subarray}}\biggl(\frac{1}{|\mathcal{A}|}\biggr)^{(2n(k-1)+\operatorname{len}(p^{\prime}))(1/2-\epsilon)}.$
Since there are $|\mathcal{A}|^{\,n-\operatorname{len}(p_{j})/2}$ choices for each $q_{j}$, the preceding display equals
$\biggl(\frac{1}{|\mathcal{A}|}\biggr)^{(2n(k-1)+\operatorname{len}(p^{\prime}))(1/2-\epsilon)}\prod_{j=1}^{k-1}|\mathcal{A}|^{\,n-\operatorname{len}(p_{j})/2}.$
A short rearrangement yields
$\sum_{U^{\prime}\in\operatorname{Extend}_{n}(U_{i})}(\operatorname{diam}U^{\prime})^{\,1/2-\epsilon}=|\mathcal{A}|^{\,\epsilon(2n(k-1)+\operatorname{len}(p^{\prime}))}\cdot\bigl(\operatorname{diam}U_{i}\bigr)^{1/2}.$
Since $\epsilon(2n(k-1)+\operatorname{len}(p^{\prime}))\leq\epsilon_{0}\operatorname{len}(p)$ and $(\operatorname{diam}U_{i})^{\epsilon_{0}}=|\mathcal{A}|^{-\epsilon_{0}\operatorname{len}(p)}$, we obtain
$\sum_{U^{\prime}\in\operatorname{Extend}_{n}(U_{i})}(\operatorname{diam}U^{\prime})^{\,1/2-\epsilon}\leq(\operatorname{diam}U_{i})^{1/2-\epsilon_{0}},$
as claimed.
It follows that
$H^{1/2-\epsilon}(W_{n})\leq\liminf_{\delta\to 0}\Biggl\{\sum_{i\in I}\sum_{U^{\prime}\in\operatorname{Extend}_{n}(U_{i})}(\operatorname{diam}U^{\prime})^{\,1/2-\epsilon}\;\Bigm|\;\{U_{i}\}_{i\in I}\ \text{is an open }\delta\text{-cover of }Z_{\mathrm{concat}}\Biggr\}$
$\leq\liminf_{\delta\to 0}\Biggl\{\sum_{i\in I}(\operatorname{diam}U_{i})^{\,1/2-\epsilon_{0}}\;\Bigm|\;\{U_{i}\}_{i\in I}\ \text{is an open }\delta\text{-cover of }Z_{\mathrm{concat}}\Biggr\}=0.$
Hence $\dim_{H}(W_{n})<1/2$.
Denote
$A:=\bigcup_{p\in Z}\operatorname{Extend}_{n}(p)\subseteq\mathcal{A}^{2n}.$
Note that
$|A|=\sum_{p\in Z}|\mathcal{A}|^{\,n-(\operatorname{len}(p)/2)}.$
Since $A_{\mathrm{concat}}\subseteq W_{n}$ and $\dim_{H}(W_{n})<1/2$, Theorem 2.19 implies $|A|<|\mathcal{A}|^{n}$. Therefore
$\sum_{p\in Z}\frac{1}{|\mathcal{A}|^{\,\operatorname{len}(p)/2}}<1,$
which is the desired inequality.
For the converse, assume
$\sum_{p\in Z}\biggl(\frac{1}{|\mathcal{A}|}\biggr)^{d\cdot\operatorname{len}(p)}<1\quad\text{for }d=\tfrac{1}{2}.$
By continuity in the exponent $d$, there exists $\epsilon>0$ such that
$\sum_{p\in Z}\biggl(\frac{1}{|\mathcal{A}|}\biggr)^{(1/2-\epsilon)\operatorname{len}(p)}<1.$
For each $n\geq 1$ the family
$\bigcup_{(p_{1},\dots,p_{n})\in Z^{n}}[T_{p_{1}\circ\cdots\circ p_{n}}]$
is an open cover of $Z_{\mathrm{concat}}$, and
$\max_{(p_{1},\dots,p_{n})\in Z^{n}}\operatorname{diam}([T_{p_{1}\circ\cdots\circ p_{n}}])=|\mathcal{A}|^{-n\cdot\min_{p\in Z}\operatorname{len}(p)}\underset{n\to\infty}{\to}0.$
Hence
$H^{1/2-\epsilon}(Z_{\mathrm{concat}})\leq\lim_{n\to\infty}\sum_{(p_{1},\dots,p_{n})\in Z^{n}}\operatorname{diam}([T_{p_{1}\circ\cdots\circ p_{n}}])^{\,1/2-\epsilon}$
$\lim_{n\to\infty}\sum_{(p_{1},\dots,p_{n})\in Z^{n}}\biggl(\frac{1}{|\mathcal{A}|}\biggr)^{(1/2-\epsilon)\sum_{j=1}^{n}\operatorname{len}(p_{j})}=\lim_{n\to\infty}\Biggl(\sum_{p\in Z}\biggl(\frac{1}{|\mathcal{A}|}\biggr)^{(1/2-\epsilon)\operatorname{len}(p)}\Biggr)^{\!n}.$
By the assumption the last expression vanishes, therefore $H^{1/2-\epsilon}(Z_{\mathrm{concat}})=0$ and $\dim_{H}(Z_{\mathrm{concat}})<1/2$.
@@ 4 Introduction to Nielsen-Schreier Theory
In Figure 1, we illustrate the Schreier graphs of two free subgroups of $F(\mathcal{A})$ with $\mathcal{A}=\{a,b\}$, one of which being the trivial subgroup, denoted $\langle\rangle$, containing only the empty word [^6] The Schreier graph of the trivial subgroup is better known as the Cayley graph of the free group..
@@ 5 Algebraic Criteria Using Maximal Prefix Codes
Notice that the proof of the previous Theorem, showed that in fact, we have at least $|\{p\in Z:\lfloor\mathrm{len}(p)/2\rfloor\text{ is maximal}\}|/|\mathcal{A}|\geq 1$ different ways to choose $Z^{\prime}$, as the initial choice of the maximal length word was arbitrary.
@@ 6 Covering the Tree by a Schreier Graph
7 Further Discussion
We show that several of the preceding results extend without difficulty to the case of a countable alphabet.
Let $\mathcal{A}$ be a countable set (finite or infinite), and let $\mu$ be a probability measure on the measurable space $(\mathcal{A},\mathcal{P}(\mathcal{A})).$ Thus $\mu$ is uniquely determined by the weights
$\mu_{a}:=\mu(\{a\}),\qquad a\in\mathcal{A},$
which satisfy
$\sum_{a\in\mathcal{A}}\mu_{a}=1.$
Remark 7.1
The assumption that the alphabet $\mathcal{A}$ is countable is essential for our arguments, as it allows the use of countable subadditivity (in particular, the union bound). It also ensures the existence of purely atomic probability measures, i.e. measures for which every singleton $\{a\}$ has positive mass $\mu_{a}>0$.
Indeed, purely atomic probability measures can only be supported on countable sets. To see this, for each $n\in\mathbb{N}$ define
$A_{n}:=\{a\in\mathcal{A}:\mu_{a}>1/n\}.$
Each set $A_{n}$ is finite, and since
$\mathcal{A}=\bigcup_{n\in\mathbb{N}}A_{n},$
it follows that $\mathcal{A}$ is countable.
Throughout this section, we consider the game $(T,W)$, where $T=\mathcal{A}^{\leq\mathbb{N}}$ and the winning set $W$ is open and of the form
$W=\bigcup_{p\in Z}[T_{p}],\qquad Z\subseteq T.$
We begin by extending Theorem 2.7 to this more general setting.
Lemma 7.2
Let $\{\mu^{i}\}_{i=1}^{\infty}$ be a sequence of probability measures on $\mathcal{A}$. If
$\sum_{p=\langle a_{1},a_{2},\dots,a_{n}\rangle\in Z}\;\prod_{i=1}^{\lfloor n/2\rfloor}\mu^{i}_{a_{2i}}<1,$
then Player $2$ has a winning strategy in the game $(T,W)$.
Proof. Assume, towards a contradiction, that Player $1$ has a winning strategy, and fix such a strategy $s_{1}^{\ast}$. Suppose that Player $2$ plays randomly as follows: at each even stage $2i$ ($i\geq 1$), Player $2$ independently selects an action $a\in\mathcal{A}$ according to the probability measure $\mu^{i}$. This induces a probability distribution on the set of Player $2$ strategies, which we denote by $s_{2}$.
Let $x(s_{1}^{\ast},s_{2})$ denote the (random) play generated by the pair of strategies $(s_{1}^{\ast},s_{2})$. Since $s_{1}^{\ast}$ is assumed to be winning for Player $1$, we have
$\mathbb{P}\bigl(x(s_{1}^{\ast},s_{2})\in W\bigr)=1.$
We recall that $W=\bigcup_{p\in Z}[T_{p}]$, and it follows that
$1=\mathbb{P}\!\left(\bigcup_{p\in Z}\{p\preceq x(s_{1}^{\ast},s_{2})\}\right)\leq\sum_{p\in Z}\mathbb{P}\bigl(p\preceq x(s_{1}^{\ast},s_{2})\bigr),$
where the inequality follows from countable subadditivity.
Now fix $p=\langle a_{1},a_{2},\dots,a_{n}\rangle\in Z$. If $p$ is not consistent with the strategy $s_{1}^{\ast}$, then $\mathbb{P}\bigl(p\preceq x(s_{1}^{\ast},s_{2})\bigr)=0$. Otherwise, the event $\{p\preceq x(s_{1}^{\ast},s_{2})\}$ occurs precisely when Player $2$ selects the actions $a_{2},a_{4},\dots,a_{2\lfloor n/2\rfloor}$ at the corresponding stages. By independence of Player $2$’s random choices, we obtain
$\mathbb{P}\bigl(p\preceq x(s_{1}^{\ast},s_{2})\bigr)=\prod_{i=1}^{\lfloor n/2\rfloor}\mu^{i}_{a_{2i}}.$
Combining the above yields
$1\leq\sum_{p=\langle a_{1},a_{2},\dots,a_{n}\rangle\in Z}\prod_{i=1}^{\lfloor n/2\rfloor}\mu^{i}_{a_{2i}},$
which contradicts the assumption of the lemma. The conclusion follows.
Remark 7.3
Let $W\subseteq[\mathcal{A}^{\leq\mathbb{N}}]$ be a winning set for which Player $1$ has a unique winning strategy. Then the sum appearing in Lemma 7.2 is equal to $1$. Indeed, in this case the events corresponding to distinct prefixes $p\in Z$ are mutually disjoint, since the probability that a given position is a prefix of two distinct plays is zero. Consequently, the inequality arising from the union bound becomes an equality.
By suitably modifying Definition 3.1 to allow a countable alphabet, one readily verifies that Lemma 3.2 and (a slightly modified) Theorem 3.5 continue to hold in this setting.
We may likewise formulate and prove an extension of Theorem 6.6 to the countable-alphabet setting.
Proposition 7.4
Let $C$ be a maximal prefix code over the alphabet $\mathcal{A}$, let $x=(x_{i})_{i\in\mathbb{N}}\in\mathcal{A}^{\mathbb{N}}$, and let $\mu$ be a probability measure on $\mathcal{A}$. Then
$\sum_{c=\langle c_{1},\ldots,c_{n}\rangle\in C}2^{\lvert\{i\in[n]:c_{i}\neq x_{i}\}\rvert}\prod_{i=1}^{n}\frac{\mu_{c_{i}}}{2-\mu_{x_{i}}}=1.$
Proof. We follow the proof of Proposition 6.6. Consider a subtree of the covering Schreier graph $\mathrm{Sch}(\langle\rangle)$ corresponding to a winning strategy of Player 1.
For each integer $i\geq 1$, the move $a_{2i-1}$ of Player $1$ at stage 2i12i-1 is uniquely determined along this subtree and is equal to either $x_{i}$ or $x_{i}^{-1}$. We define a family of probability measures $\{\mu^{i}\}_{i\geq 1}$ on $\mathcal{A}\cup\mathcal{A}^{-1}\setminus\{a_{2i-1}^{-1}\}$ as follows. For every $a\in\mathcal{A}\setminus\{x_{i}\}$, set
$\mu^{i}_{a}=\mu^{i}_{a^{-1}}:=\frac{\mu_{a}}{2}\cdot\frac{1}{1-(\mu_{x_{i}}/2)}=\frac{\mu_{a}}{2-\mu_{x_{i}}}.$
Also set $\mu^{i}_{a_{2i-1}}=\frac{\mu_{x_{i}}}{2-\mu_{x_{i}}}$. Each $\mu^{i}$ is well defined and satisfies
$\sum_{a\in\mathcal{A}\cup\mathcal{A}^{-1}\setminus\{a_{2i-1}^{-1}\}}\mu^{i}_{a}=\frac{\mu_{x_{i}}}{2-\mu_{x_{i}}}+2\sum_{a\in\mathcal{A}\setminus\{x_{i}\}}\frac{\mu_{a}}{2-\mu_{x_{i}}}=\frac{\mu_{x_{i}}}{2-\mu_{x_{i}}}+2\cdot\frac{1-\mu_{x_{i}}}{2-\mu_{x_{i}}}=1.$
Applying the argument used in the proof of Theorem 6.6, together with Lemma 7.2, to the measures $\{\mu^{i}\}_{i\geq 1}$ and the subtree determined by the fixed strategy of Player $1$, yields the desired conclusion.
Remark 7.5
For simplicity, the probability measures assigned at each “level” of the Schreier graph were chosen to be symmetric, assigning equal mass to the singleton sets $\{a\}$ and $\{a^{-1}\}$. This symmetry assumption is not necessary and the measures could be modified freely. However, doing so would substantially complicate the resulting expressions.
Example 7.6
Let $\mathcal{A}=\mathbb{N}_{>0}=\{1,2,\dots\}$, and let $\mu$ be a probability measure on $\mathcal{A}$ defined by
$\mu(\{n\})=a_{n},\qquad n\in\mathbb{N}_{>0},$
where $(a_{n})_{n\geq 1}$ is a sequence of non-negative real numbers satisfying
$\sum_{n=1}^{\infty}a_{n}=1.$
We illustrate the results with the choice $a_{n}=2^{-n}$. For every maximal prefix code $C\subseteq\mathcal{A}^{<\mathbb{N}}$,
$\sum_{c=\langle c_{1},\ldots,c_{n}\rangle\in C}2^{-\sum_{i=1}^{n}c_{i}}=1.$
Moreover, for every sequence $x=(x_{i})_{i\in\mathbb{N}}\in\mathcal{A}^{\mathbb{N}}$,
$\sum_{c=\langle c_{1},\ldots,c_{n}\rangle\in C}2^{\sum_{i=1}^{n}(x_{i}-c_{i}+1)\,\mathbf{1}_{\{x_{i}\neq c_{i}\}}}\prod_{i=1}^{n}\frac{1}{2^{x_{i}+1}-1}=1.$
Example 7.7
Let $\mathcal{A}$ be a finite alphabet, and define the uniform probability measure
$\mu_{a}=\frac{1}{|\mathcal{A}|},\qquad a\in\mathcal{A}.$
In this case, the results of the present section reduce to the main results established in Section 2.